OpenStudy (anonymous):

* I'll answer your questions pertaining to English, Psych, and Philosophy if you help me with mine? (My answer is D) Solve A = 6p - 2r for p. p = A - 2r p = A + 4r p = a-2r/-6 p = A+2r/6

OpenStudy (anonymous):

Need to get p on its own again, so dealing with the right hand side we want to get rid of the -2r. We do this by adding 2r to both sides, giving: \[A+2r=6p\] Now we have 6p, but we want just p. So we divide both sides by 6 to get: \[\frac{A+2r}{6}=p\]

OpenStudy (anonymous):

You add 2r to 6p?

OpenStudy (anonymous):

Ok, thank you!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

I added 2r to \(6p-2r\), because that gives \(6p-2r+2r\) and the 2r's cancel :)

OpenStudy (anonymous):

The best way to think about these kind of questions is 'unravelling' the equation to find what you want. We want to find p. The equation says that if we multiply this number by 6, then subtract 2r from it, we get A. We need to unravel these operations, so: -First we should ADD 2r (opposite of subtracting 2r). -Then we should divide by 6 (opposite of multiplying by 6). Does that help?

OpenStudy (anonymous):

The 2r's cancel---- got it

OpenStudy (anonymous):

You're so smart, I'm jealous...

OpenStudy (anonymous):

Great :) Now in response I want to know the meaning of life!

OpenStudy (anonymous):

Ha thanks, you're being smart by coming on here to study :)

OpenStudy (anonymous):

I'll make it brief, yes?

OpenStudy (anonymous):

Oh go on then.

OpenStudy (anonymous):

That's one way of looking at it, thank you.

OpenStudy (anonymous):

Give me a second while I collect my thoughts.

OpenStudy (anonymous):

This is just a glimpse of how I think: When you think about it, everyone is struggling. Whatever you may be going through possibly seems like the toughest mechanism, but in comparison to what others are going through, they may see it as the meekest of complications. I don’t necessarily think that anytime you face hardships that you should compare it to major problems that others are encountering, but I do think you should be aware of no matter what the situation is, someone in the world is or are going through it right now. I know what it is like to not want to focus on positive things. But when you are in a place of peace, you can recognize the kinds of things you never noticed before. Give yourself the opportunity to see the beauty in the world, and if you’re having trouble with that, keep on looking because there is magnificence in the world.

OpenStudy (anonymous):

I didn't expect that much :P

OpenStudy (anonymous):

I told you I have a lot to offer when it comes to those sort of subjects

OpenStudy (anonymous):

Math is not my strong suit whatsoever, and never will be, but psychology and philosophy will always be something I'll be able to contribute to. As well as English.

OpenStudy (anonymous):

Excellent :) Everybody has their strengths. Is there a particular philosopher you particularly enjoy reading?

OpenStudy (anonymous):

I didn't mean to use particular twice, but I'm allowed because it's late.

OpenStudy (anonymous):

No worries whatsoever.

OpenStudy (anonymous):

And yes, Aristotle is a favorite

OpenStudy (anonymous):

Oh great :D Have you read any Plato or Socrates?

OpenStudy (anonymous):

Some of Plato, not much of Socrates

OpenStudy (anonymous):

I'm very into self-improvement books

OpenStudy (anonymous):

And any book covering Psychology is a weakness of mine

OpenStudy (anonymous):

Ah right. Have you heard about Ayn Rand?

OpenStudy (anonymous):

Not per say, no

OpenStudy (anonymous):

However, he's Russian and American, as well as I

OpenStudy (anonymous):

She's pretty cool, all about individualism.

OpenStudy (anonymous):

she's***

OpenStudy (anonymous):

My error

OpenStudy (anonymous):

Thank you for your help

OpenStudy (anonymous):

No problem :)